OpenStudy (anonymous):

Two sides of a triangle have lengths 8 and 17. What must be true about the third side?

OpenStudy (anonymous):

That the length of the the third side must be <8+17=25

OpenStudy (anonymous):

I dont understand

hero (hero):

If you have a triangle with sides a, b, and c and you know the lengths of sides a and b, then the length of c will be: |a - b| < c < |a + b|

hero (hero):

In other words, in this case a = 8 b = 17 So c is |8 - 17| < c < |8 + 17|

hero (hero):

@BBeckaBee what is the range of length c? Did you figure it out?

hero (hero):

Do you know what absolute value symbols || mean?

hero (hero):

|8 - 17| = |-9| = 9 The absolute value of a negative number is a positive number.

hero (hero):

And |8 + 17| = |25| = 25

hero (hero):

So this means that the length of c will be 9 < c < 25

hero (hero):

In other words, the length of c will be between 9 and 25 given that the length of a is 8 and the length of b is 17

hero (hero):

This is standard procedure for figuring out the range of the length of side c.

hero (hero):

@BBeckaBee I'm sorry you don't understand.